how many ml of a 4% saline solution must be mixed with 40ml of a 10% saline solution to get a 6% saline solution?
Answer by mananth(16949)

how many ml of a 4% saline solution must be mixed with 40ml of a 10% saline solution to get a 6% saline solution?

4% saline solution -------------x ml
10% saline solution-------------40 ml
6% saline solution-------------(x+40)ml
..
4x+10*40= 6(x+40)
4x+400=6x+240
6x-4x=400-240
2x=160
/2
x=80 ml. of 6% solution